OpenStudy (anonymous):

Find dx/dy for the following function: y=sinx+5e^.4x dx/dy =

OpenStudy (anonymous):

is it dx/dy?

OpenStudy (anonymous):

what does dx/dy equals

OpenStudy (anonymous):

I just want to make sure the question, because dx/ dy !=dy/dx

OpenStudy (anonymous):

if it is dx/ dy. it's not easy to solve. you must solve for x first , then take derivative respect to y.

OpenStudy (anonymous):

it will behave implicitly

OpenStudy (anonymous):

i think she means dy/dx

OpenStudy (anonymous):

i thought i had to find the derivative and then solve for y

OpenStudy (anonymous):

@mathsmind you help her, ok?

OpenStudy (anonymous):

no u help her

OpenStudy (anonymous):

can u double check the question plz

OpenStudy (anonymous):

it must be dy/dx at ur level

OpenStudy (anonymous):

\[Find \frac{ dx }{ dy } for the following function. Y = sinx + 5e ^{0.4x}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

u allowed everyone to escape from this question hehehehe

OpenStudy (anonymous):

lol i know, i wish i could escape it too

OpenStudy (anonymous):

dx/dy = 1/(dy/dx)

OpenStudy (anonymous):

@jim_thompson5910 rescue me please

OpenStudy (anonymous):

ok i will solve it all it is implicit differentiation but in terms of x

OpenStudy (anonymous):

so when u diff sin(x) it will be cos(x)dx/dy

OpenStudy (anonymous):

and y will become 1 when u differentiate it

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Just find dy/dx and take the reciprocal.

OpenStudy (anonymous):

nope don't do that

OpenStudy (anonymous):

they are not the same

OpenStudy (anonymous):

Look, dx/dy = 1/(dy/dx)

OpenStudy (anonymous):

no it does not work like that

OpenStudy (anonymous):

I hope you're kidding. I'm not manipulating fractions

OpenStudy (anonymous):

look multiply both sides

OpenStudy (anonymous):

u r missing the chain rule this is calculus not algebra

OpenStudy (anonymous):

the case where u use a reciprocal is by using the chain rule and cancelling each term out

OpenStudy (anonymous):

Can't you implicit differentiate in this situation?

OpenStudy (anonymous):

implicitly*

OpenStudy (anonymous):

yes that is different from ur fomula

OpenStudy (anonymous):

im even more confused...

OpenStudy (anonymous):

he wants u to use L rule

OpenStudy (anonymous):

which is similar to what i said about the chain rule

OpenStudy (anonymous):

Lebenz derived a formula from the chain rule

OpenStudy (anonymous):

\[\large y=\sin x+5e^{0.4x}\] \[\large 1=(\cos x +2e^{0.4x})\frac{dx}{dy}\]

OpenStudy (anonymous):

nope i told u don't do that, unless if u want to apply Lebenz rule u need to take the inverse of the function then take the reciprocal ...

OpenStudy (anonymous):

because calculus is about functions and expansions

OpenStudy (anonymous):

read that carefully If the derivative of y = f(x) is dy/dx, then the derivative of the inverse function which expresses x in terms of y is given by the formula

OpenStudy (anonymous):

in other words ur Lebenz works if x= sin(y)+5e^y

OpenStudy (anonymous):

so u can take the inverse of the function then apply the rule ok

OpenStudy (anonymous):

i told u they are not the same but thanks anyway for bribing the topic up

OpenStudy (anonymous):

so take the inverse of the function then use the rule

OpenStudy (anonymous):

or solve it implicitly as i started

OpenStudy (anonymous):

if the case was simple like this then implicit differentiation would have never exist

OpenStudy (anonymous):

\[y = \sin(x) + e^{0.4x} \longrightarrow 1=\frac{dx}{dy}\cos(x)+2\frac{dx}{dy}e^{0.4x}\]

OpenStudy (anonymous):

the method u use is solving this in an implicit way

OpenStudy (anonymous):

take dx/dy as a common factor

OpenStudy (anonymous):

well @Xavier ur method also works congratulation!

OpenStudy (anonymous):

so both methods work fine, actually u made me come with a new theory in math, or a different proof for Leben'z formula. although maths is not my major, but i will write a new paper on this and publish it

OpenStudy (anonymous):

works*

OpenStudy (anonymous):

not just that i will change my proof for Kepler's laws of motion ...